The table summarizes results from 975 pedestrian deaths that were caused by automobile accidents.

Pedestrian Deaths
Driver
Intoxicated?
Pedestrian Intoxicated?
Yes No
Yes 41 68
No 228 638

If one of the pedestrian deaths is randomly selected, find the probability that the pedestrian was intoxicated or the driver was not intoxicated. Please enter a decimal to 4 decimal places.

Probability = _____
Answer #1

P( the pedestrian was intoxicated or the driver was not intoxicated)=P(the pedestrian was intoxicated)+P( driver was not intoxicated)-P(the pedestrian was intoxicated and the driver was not intoxicated)

P(the pedestrian was intoxicated)=(number of deaths due to intoxicated pedestrians)/975=(41+228)/975

P( driver was not intoxicated)=(number of deaths due to not toxicated drivers)/975=(228+638)/975

P(the pedestrian was intoxicated and the driver was not intoxicated)=(number of deaths due to intoxicated pedestrians and non toxicated drivers)/975=228/975

Thus P( the pedestrian was intoxicated or the driver was not intoxicated)=(269+866-228)/975=0.9302564